A diver observes a bubble of air rising from the bottom of a lake (where the absolute pressure is 3.45 atm) to the surface (where the pressure is 1.00 atm). The temperature at the bottom is 3.9°C, and the temperature at the surface is 21.3°C.

(a) What is the ratio of the volume of the bubble as it reaches the surface to its volume at the bottom?
Answer #1

\frac{p_{1}V_{1}}{T_{1}}=\frac{p_{2}V_{2}}{T_{2}}

1=parameters at the bottom of the lake

2=parameters at the surface

T1=276.9 K

T2=294.3 K

Replace the numbers and get: V2/V1=3.667.